# Reconstructing Satellites in 3D from Amateur Telescope Images

### v3

@@ Abstract
Monitoring space objects is crucial for space situational awareness, yet reconstructing 3D satellite models from ground-based telescope images is super challenging due to atmospheric turbulence, long observation distances, limited viewpoints, and low signal-to-noise ratios. In this paper, we propose a novel computational imaging framework that overcomes these obstacles by integrating a hybrid image pre-processing pipeline with a joint pose estimation and 3D reconstruction module based on controlled Gaussian Splatting (GS) and Branch-and-Bound (BnB) search. We validate our approach on both synthetic satellite datasets and on-sky observations of China’s Tiangong Space Station and the International Space Station, achieving robust 3D reconstructions of low-Earth orbit satellites from ground-based data. Quantitative evaluations using SSIM, PSNR, LPIPS, and Chamfer Distance demonstrate that our method outperforms state-of-the-art NeRF-based approaches, and ablation studies confirm the critical role of each component. Our framework enables high-fidelity 3D satellite monitoring from Earth, offering a cost-effective alternative for space situational awareness. Project page: https://ai4scientificimaging.org/ReconstructingSatellites.
@@ 1 Introduction
However, 3D reconstruction has been only investigated in the space-based space surveillance missions, i.e. leveraging the videos captured by satellites to retrieve the 3D information of its neighboring space objects. 3D reconstruction from ground-based observations, despite its advantages in terms of cost-effectiveness and versatility, remains under-explored because of a series of unique challenges compared with its space-based counterparts, including but not limited to: 1) atmospheric turbulence significantly distorts images captured by ground telescopes; 2) the extended focal length of the ground telescope complicates the accurate estimation of a satellite’s pose; and 3) the rapid motion of satellites necessitates short exposure times, resulting in low signal-to-noise (SNR) ratios and constrained observational perspectives. As a result, classical 3D reconstruction approaches, no matter geometry or learning-based, fail in this challenging scenario.
@@ 2.2 Imaging through Turbulence
Neural Radiance Fields (NeRF) [34, 35] have revolutionized 3D reconstruction and view synthesis by modeling a scene’s volumetric density and color using deep networks. This implicit representation enables photorealistic renderings that capture intricate lighting and transparency from sparse viewpoints[36, 37, 38]. Variants such as NeRF--- - and BARF further incorporate motion information inspired by SfM, allowing simultaneous 3D reconstruction and pose estimation from unposed images. However, despite their ability to capture fine geometric and appearance details, NeRF-based methods typically require significant training time and offer limited ease for scene editing or post-processing.
@@ 3 Method
Our method addresses the challenging problem of 3D reconstruction of space objects from unposed, monocular images captured by ground-based telescopes. In contrast to conventional 3D reconstruction tasks, our approach must contend with several unique challenges: atmospheric turbulence, extremely long observation distance, limited viewpoints, and sparse, low-quality features. As shown in Fig.2 Our method overcomes these issues through two key innovations:1) a hybrid image pre-processing pipeline combining classical and deep learning techniques, and 2) a joint pose-refinement and 3D reconstruction framework using Gaussian Splatting (GS) with controlled, iterative optimization.
@@ 3.1.1 Turbulence Mitigation via Lucky Imaging
We begin by applying lucky imaging [46] to high-frame-rate, short-exposure image sequences, statistically selecting the top 12% of frames based on sharpness and contrast metrics. These selected frames then undergo centroid-based alignment to correct for bulk satellite motion, followed by intensity-normalized cross-correlation to address residual sub-pixel jitter. Finally, the well-aligned frames are stacked to improve the signal-to-noise ratio.
@@ 3.1.2 Multi-Scale Sharpening with Wavelet Decomposition
While stacking reduces noise, it may also blur high-frequency details critical for reconstruction. To counteract these effects, we apply Gaussian wavelet decomposition to separate the image into different frequency bands, selectively amplifying high-frequency details (e.g., satellite edges and surface textures) while suppressing low-frequency background light. This effectively improve the clarity and contrast of the ground telescope images, making the features of space objects more distinguishable.
@@ 3.2.1 Initialization Using Modified SfM
Our approach first leverages pre-trained deep learning models—SuperPoint for feature extraction and SuperGlue for feature matching—to detect and pair keypoints robustly despite distortions, blur, and sparse textures inherent in telescope images. To mitigate the challenges posed by the long observation distance, we adopt an orthographic projection model (instead of a perspective model), which is more robust when perspective effects are minimal. Applying this model to the first three frames yields an initial point cloud along with corresponding camera poses.
Next, we sequentially incorporate additional frames into our model, as shown in the ”Modified SfM” module in Fig.2 (b). For each new image, we first apply the Random Sample Consensus (RANSAC) algorithm to filter out incorrect feature correspondences between adjacent images, ensuring more accurate feature matches and better associations with the 3D point cloud. Based on these refined correspondences, we then estimate the initial camera pose using the Perspective-n-Point (PnP) algorithm. Additionally, to prevent error accumulation, we perform a global bundle adjustment after each new frame is added, refining all camera poses simultaneously. Additionally, every five frames we perform triangulation and integrate new features (e.g., from the 5th and 6th frames, then the 10th and 11th, etc.) into the point cloud. This iterative process continues until all images have been processed, resulting in a complete initial 3D model and the satellite’s pose estimates.
@@ 3.2.2 Controlled Gaussian Splatting with BnB Pose Refinement
The core of our reconstruction pipeline is a novel integration of 3D Gaussian splatting with iterative pose refinement. We begin by initializing Gaussian points from the SfM-derived point cloud, then iteratively optimize the GS parameters while holding camera poses fixed, and vice versa. This alternating optimization enables simultaneous recovery of the satellite’s 3D shape and accurate camera poses despite challenges such as long observation distance, limited viewpoints, and sparse features.
In early GS training phase, we fix the number of Gaussians and update only their parameters. At early stages, the limited number of Gaussians forces each one to cover a larger area, effectively capturing the satellite’s rough geometry while avoiding overfitting to high-frequency artifacts when poses are imprecise. After a set number of epochs, we automatically clone and split the Gaussians to increase their count, then resume training on the expanded set with the count held fixed. This controlled growth strategy enables the model to first learn coarse geometry and gradually capture finer details.
In the first 10,000 epochs of training, we refine the camera poses through a Branch-and-Bound (BnB) search every 1,000 epochs. Candidate poses are generated by applying controlled perturbations to the current estimate. Specifically, for rotation, we sample a random rotation angle, $\Delta\theta$, and a random axis, $\vec{n}$, then compute the rotation perturbation using Rodrigues’ formula:
@@
where $[\vec{n}]_{\times}$ is the skew-symmetric matrix of $\vec{n}$. Translation perturbations are drawn from a zero-mean Gaussian distribution:
@@
The newly generated camera poses are shown in red in Fig.2 (b). We evaluate these candidates using a photometric consistency loss and select the pose $(R_{\text{best}},T_{\text{best}})$ that minimizes this loss. In our scenario, where long observation distance and limited viewpoints make pose estimation highly sensitive, the BnB search constrains the adjustment range, ensuring robust optimization even in the presence of significant noise. The refined pose is then fed back into the next GS training stage. By iterating between controlled Gaussian splatting with adaptive point growth and BnB-based pose refinement over roughly ten cycles, our method progressively converges to an optimal joint reconstruction of the satellite’s 3D model and its pose.
@@ 3.2.3 Explicit Post-processing of the Learned Point Cloud
The reconstructed GS point cloud may still contain stray points from background noise. To enhance the 3D model’s accuracy, we apply explicit post-processing using statistical filtering to remove outlier Gaussian points. As shown in Fig.2 (c), for each point, we compute the average distance to its $k$ nearest neighbors and prune those whose mean distance exceeds the overall mean plus one standard deviation ($\mu+\sigma$). After filtering, we resume Gaussian splatting training on the refined point cloud for an additional 500 epochs to repair discontinuities and further improve reconstruction quality.
@@ 4.1.1 Synthetic Ground Telescope Satellite Images
We generated three distinct 3D satellite models using Blender, each approximately 60 meters in size, and simulated ground telescope images for each. Placed in a low-Earth orbit at 650 km altitude, each satellite is visible from a stationary ground telescope in a 13 segment of a full 360 trajectory. Our virtual telescope—with a 0.35-meter aperture and a 3.2-meter focal length—rendered 700 high-quality images per model. To simulate realistic observing conditions, each clean image was augmented with 20 different atmospheric turbulence settings (Fried parameter between 0.07 and 0.35 m) and 5–7% additional background intensity due to light pollution. This resulted in a dataset of 14,000 distorted frames per model, each with a resolution of 512 $\times$ 512 pixels, which we will publicly release after the paper is accepted.
@@ 4.1.2 Raw Telescope Image Pre-processing
We first validate our pre-processing pipeline on the simulated data. Following our pipeline—lucky imaging and stacking (every 100 images), wavelet-based multi-scale sharpening, and VRT denoising—we reduce the 14,000 distorted frames to 140 high-quality images per satellite model. We compared our method against other cutting-edge deblurring algorithms, Table I summarizes the evaluation metrics (SSIM, PSNR, LPIPS) across various preprocessing methods. The results indicate that our proposed pipeline (Stacking + Sharpening + VRT) consistently surpasses pure deep learning approaches (e.g., Restormer[51]), traditional techniques (e.g., Stacking + Sharpening), self-supervised methods, and other hybrid strategies, specifically in PSNR and SSIM metrics.
@@ 4.1.3 Robust Pose Initialization
From the 140 processed images, we select every tenth frame (e.g., 1st, 10th, 20th) for the training set and reserve the intermediate frames for validating the reconstruction. We first compare the pose estimation accuracy of our method against existing techniques, including feature-based SfM methods—such as Colmap[31], hierarchical localization[52], and detector-free SfM[53]—as well as several feed-forward deep learning approaches, including DUSt3R[54] and RayDiffusion[55]. Figure 4 and Table II compare the estimated poses to the ground truth (GT) on the training set. Feature-based SfM methods accumulate noise, which leads to inaccurate keypoint matching and causes bundle adjustment (BA) to fail to converge; consequently, no reasonable pose estimation can be reported. Similarly, deep learning methods struggle with extreme focal lengths, structural symmetry, and poor textures, resulting in incorrect depth estimation and subsequent pose errors. Other approaches, such as PoseDiffusion[56], MV-DUSt3R+[57], and Colmap-free 3DGS[58], also exhibit substantial pose estimation errors. In contrast, our pose initialization method achieves an ATE of only 2.61, with the estimated pose distribution closely aligning with the ground truth. As explained in Sec.3.2.1, our superior results stem from three key factors: (1) a robust orthographic projection model; (2) the use of RANSAC algorithms to eliminate erroneous matches; and (3) a strategy that leverages only the matching relationships between adjacent frames to relax the convergence conditions. These design choices enable our method to overcome the limitations of traditional SfM and deep learning approaches, significantly enhancing reliability for pose estimation of texture-poor space objects.

### v5

@@ Abstract
Monitoring space objects is crucial for space situational awareness, yet reconstructing 3D satellite models from ground-based telescope images is super challenging due to atmospheric turbulence, long observation distances, limited viewpoints, and low signal-to-noise ratios. In this paper, we propose a novel computational imaging framework that overcomes these obstacles by integrating a hybrid image pre-processing pipeline with a joint pose estimation and 3D reconstruction module based on controlled Gaussian Splatting (GS) and Branch-and-Bound (BnB) search. We validate our approach on both synthetic satellite datasets and on-sky observations of China’s Tiangong Space Station and the International Space Station, achieving robust 3D reconstructions of low-Earth orbit satellites from ground-based data. Quantitative evaluations using SSIM, PSNR, LPIPS, and Chamfer Distance demonstrate that our method outperforms state-of-the-art NeRF-based approaches, and ablation studies confirm the critical role of each component. Our framework enables high-fidelity 3D satellite monitoring from Earth, offering a cost-effective alternative for space situational awareness. Project page: https://ai4scientificimaging.org/3DSatellites.
@@ 1 Introduction
To date, 3D reconstruction has only been investigated in the space-based space surveillance missions, i.e. leveraging the videos captured by a satellite to retrieve the 3D information of its neighboring space objects. 3D reconstruction from ground-based observations, despite its advantages in terms of cost-effectiveness and versatility, remains under-explored because of a series of unique challenges compared with its space-based counterparts, including but not limited to: 1) atmospheric turbulence significantly distorts images captured by ground telescopes; 2) the long observation distance complicates the accurate estimation of a satellite’s pose; and 3) the rapid motion of satellites necessitates short exposure times, resulting in low signal-to-noise (SNR) ratios and constrained observational perspectives. As a result, classical 3D reconstruction approaches, no matter geometry or learning-based, fail in this challenging scenario.
@@ 2.2 Imaging through Turbulence
Neural Radiance Fields (NeRF) [34, 35] have revolutionized 3D reconstruction and view synthesis by modeling a scene’s volumetric density and color using deep networks. This implicit representation enables photorealistic renderings that capture intricate lighting and transparency from sparse viewpoints[36, 37]. Variants such as NeRF-- and BARF further incorporate motion information inspired by SfM, allowing simultaneous 3D reconstruction and pose estimation from unposed images. However, despite their ability to capture fine geometric and appearance details, NeRF-based methods typically require significant training time and offer limited ease for scene editing or post-processing.
@@
2.4 Equatorial Mount and Meridian Flip
In astronomical imaging, telescopes are commonly mounted on equatorial mounts [45] to compensate for the Earth’s rotation and the apparent motion of astronomical objects. These mounts have two orthogonal axes to enable smooth tracking: the right ascension (RA) axis, aligned with Earth’s rotational axis and pointing toward the celestial north pole, and the declination (Dec) axis, which adjusts the telescope’s elevation (see Fig. 2(a)). During tracking, the RA axis rotates at either the sidereal rate or the satellite’s orbital rate to keep the target centered, while the Dec axis typically requires only minor adjustments.
However, as the target crosses the local meridian—a vertical plane passing through the zenith—the RA axis must continue rotating (Fig.2(b)). This motion can bring the telescope dangerously close to the mount or tripod, risking mechanical collision (Fig.2(c)). To prevent this, most equatorial mounts execute a meridian flip [46, 47], rotating the RA axis by 180 to reposition the telescope on the opposite side (Fig. 2(d)). This operation introduces a brief interruption and a mirrored camera orientation, typically causing a pause of around 30 seconds.
While such delays are negligible for slowly drifting targets like stars or nebulae, they become problematic for fast-moving satellites, which remain visible for only 2–3 minutes per pass. Compounding this issue, the meridian region often offers the best imaging conditions due to minimal atmospheric distortion. To address these challenges in ground-based satellite tracking, we should reconfigure the mount alignment to avoid meridian flips—an approach we detail in Sec. 4.2.1, referred to as the ”Fake Polar Axis” method.
Figure 2: Illustration of equatorial mount tracking and meridian flip. (a) The telescope tracks a target across the celestial sphere. (b) The target reaches the local meridian. (c) Without adjustment, the telescope risks mechanical collision with the mount. (d) A meridian flip is performed, rotating the RA axis by 180 to safely continue tracking.
@@ 3 Method
Our method addresses the challenging problem of 3D reconstruction of space objects from unposed, monocular images captured by ground-based telescopes. In contrast to conventional 3D reconstruction tasks, our approach must contend with several unique challenges: atmospheric turbulence, extremely long observation distance, limited viewpoints, and sparse, low-quality features. As shown in Fig. 3, our method overcomes these issues through two key innovations:1) a hybrid image pre-processing pipeline combining classical and deep learning techniques, and 2) a joint pose-refinement and 3D reconstruction framework using Gaussian Splatting (GS) with controlled, iterative optimization.
@@ 3.1.1 Turbulence Mitigation via Lucky Imaging
We begin by applying lucky imaging [48] to high-frame-rate, short-exposure image sequences, statistically selecting the top 12% of frames based on sharpness and contrast metrics. These selected frames then undergo centroid-based alignment to correct for bulk satellite motion. Each frame is subsequently divided into multiple patches, and intensity-normalized cross-correlation is applied per patch to correct residual sub-pixel jitter induced by local turbulence. The aligned frames $\{\widetilde{I}_{t}\}_{t\in\mathcal{S}}$ are stacked to suppress noise and geometric distortions. The stacked image is computed as:
$I_{\text{stacked}}(x,y)=\frac{1}{|\mathcal{S}|}\sum_{t\in\mathcal{S}}\widetilde{I}_{t}(x,y),$
where $\mathcal{S}$ denotes the subset of selected sharp frames. This averaging reduces zero-mean random noise by a factor of $1/\sqrt{|\mathcal{S}|}$ and improves the signal-to-noise ratio (SNR).
@@ 3.1.2 Multi-Scale Sharpening with Wavelet Decomposition
While stacking reduces noise, it may also blur high-frequency details and suppress low-contrast structures due to short exposure. To restore such details, we apply multi-scale sharpening based on Gaussian wavelet decomposition, which separates the image into different frequency bands and selectively enhances the high-frequency components. Let $W_{j}(I)$ denote the wavelet decomposition of image $I$ at scale $j$ ($j=1,\dots,6$, from fine, $256\times 256$, to coarse, $8\times 8$). We apply a multiplicative gain factor $\alpha_{j}$ to each level:
$\widetilde{W}_{j}(I)=\alpha_{j}\cdot W_{j}(I),$
with
$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6})=(1.1,1.1,1.1,1.1,1.1,5.0).$
The substantially larger gain at level 6 compensates for global contrast loss and enhances the visibility of extended, low-contrast structures (e.g., solar wings).
@@ 3.2.1 Initialization Using Modified SfM
Our approach first leverages pre-trained deep learning models—SuperPoint for feature extraction and SuperGlue for feature matching—to detect and pair keypoints robustly despite distortions, blur, and sparse textures inherent in telescope images. To mitigate the challenges posed by the long observation distance, we adopt an orthographic projection model[50] (instead of a perspective model) in the initial step, which is more robust when perspective effects are minimal. Applying this model to the first three frames yields an initial point cloud along with corresponding camera poses.
Next, we revert to the full perspective camera model and sequentially integrate additional frames, as shown in the ”Modified SfM” module in Fig. 3 (b). For each new image, we first apply the Random Sample Consensus (RANSAC) algorithm to filter out incorrect feature correspondences between adjacent images, ensuring more accurate feature matches and better associations with the 3D point cloud. Based on these refined correspondences, we then estimate the initial camera pose using the Perspective-n-Point (PnP) algorithm. Additionally, to prevent error accumulation, we perform a global bundle adjustment after each new frame is added, refining all camera poses simultaneously. Additionally, every five frames we perform triangulation and integrate new features (e.g., from the 5th and 6th frames, then the 10th and 11th, etc.) into the point cloud. This iterative process continues until all images have been processed, resulting in a complete initial 3D model and the satellite’s pose estimates.
@@ 3.2.2 Controlled Gaussian Splatting with BnB Pose Refinement
The core of our reconstruction pipeline is a novel integration of 3D Gaussian splatting with iterative pose refinement [51]. We begin by initializing Gaussian points from the SfM-derived point cloud, then iteratively optimize the GS parameters while holding camera poses fixed, and vice versa. This alternating optimization enables simultaneous recovery of the satellite’s 3D shape and accurate camera poses despite challenges such as long observation distance, limited viewpoints, and sparse features.
During the early GS training phase, we fix the number of Gaussians and optimize only their parameters—position, covariance, opacity, and spherical harmonic coefficients (up to third order). By limiting the number of Gaussians, we force each one to cover a larger region, effectively capturing the satellite’s coarse geometry while preventing overfitting to high-frequency artifacts when pose estimates are imprecise. After a set number of epochs, we automatically clone and split the Gaussians to increase their count, then resume training on the expanded set with the count held fixed. This controlled growth strategy enables the model to first learn coarse geometry and gradually capture finer details.
In the first 10,000 epochs of training, we refine the camera poses through a Branch-and-Bound (BnB) search performed every 1,000 epochs. Candidate poses are generated by applying controlled perturbations to the current estimate. Specifically, for rotation, we sample a random rotation angle, $\Delta\theta$, and a random axis, $\vec{n}$, then compute the rotation perturbation using Rodrigues’ formula:
@@
where $\vec{n}]_{\times}$ is the skew-symmetric matrix of $\vec{n}$. Translation perturbations are drawn from a zero-mean Gaussian distribution:
@@
At each BnB stage, we sample 200 pose candidates (example camera poses are shown in red in Fig.  3(b)). We evaluate these candidates using a photometric consistency loss and select the pose $(R_{\text{best}},T_{\text{best}})$ that minimizes this loss. After each stage, we shrink the translation range to one-fourth and the rotation range to one-half of the previous interval, centering both on the current camera pose estimate. This refinement schedule ensures stable convergence within ten rounds. The best pose is then passed to the next GS training phase. By alternating between Gaussian splatting with adaptive point growth and BnB-based pose refinement, our method jointly converges to accurate 3D structure and camera poses.
Our implementation builds on the original 3D Gaussian Splatting framework [38], using the same hyper-parameter set across all experiments—synthetic data, CSS, and ISS—without scene-specific tuning, and keeping all other settings unchanged except for the modifications described above.
@@ 3.2.3 Explicit Post-processing of the Learned Point Cloud
Because the reconstructed 3D Gaussians may include stray points from background noise, we further apply a statistical filtering step to improve model accuracy. As illustrated in Fig. 3(c), we examine each Gaussian’s spatial location (ignoring its appearance and shape attributes), compute the mean distance to its $k$ nearest neighbors, and remove any point whose mean distance exceeds the global mean plus one standard deviation ($\mu+\sigma$). After filtering, we resume Gaussian splatting training on the refined 3D Gaussians for an additional 500 epochs to repair discontinuities and further improve reconstruction quality.
@@ 4 Experiments
Figure 7: Satellite tracking geometry and the Fake Polar method: (a) A 13° ground track corresponds to a 116° apparent angular sweep from the telescope. (b) Tilting the mount’s polar axis into the orbital plane avoids meridian flip and shifts tracking motion primarily to the Dec axis.
@@ 4.1.1 Synthetic Ground Telescope Satellite Images
We generated three distinct 3D satellite models using Blender, each approximately 60 meters in size, and simulated ground telescope images for each. Placed in a low-Earth orbit at 650 km altitude, each satellite is visible from a stationary ground telescope in a 13 segment, which corresponds to an apparent 116 relative rotation between the satellite and the telescope (see Fig. 7 (a)). Our virtual telescope—with a 0.35-meter aperture and a 3.2-meter focal length—rendered 700 high-quality images per model. To simulate realistic observing conditions, we introduced a full degradation model incorporating both atmospheric turbulence and light pollution. Specifically, each clean image was degraded as:
$I_{\text{noisy}}=(I_{\text{clean}}\ast\text{PSF}_{\text{turb}})+B+n_{\text{ph}}+n_{\text{read}}.$
Here, $\text{PSF}_{\text{turb}}$ is the optical point spread function degraded by atmospheric turbulence, simulated using the Phase-to-Space Transform (P2S) method [25]. P2S approximates Kolmogorov turbulence by sampling Zernike polynomials and PCA bases learned from real turbulence data. The turbulence strength is controlled via the Fried parameter $r_{0}$, which we sampled uniformly from the range [0.07m, 0.35m], equivalent to restricting the turbulence integration height to roughly the lowest 10 km of the atmosphere. Light pollution was modeled as a combination of additive background bias $B\sim\mathcal{U}(0.05\,I_{\text{clean}}^{\text{max}},0.07\,I_{\text{clean}}^{\text{max}})$, signal-dependent photon noise $n_{\text{ph}}\sim\mathcal{N}(0,\sigma_{1}^{2}(I_{\text{turb}}+B)^{2})$, and signal-independent readout noise $n_{\text{read}}\sim\mathcal{N}(0,\sigma_{2}^{2})$, where $I_{\text{turb}}=I_{\text{clean}}\ast\text{PSF}_{\text{turb}}$. This procedure resulted in a dataset of 14,000 degraded frames per model, each at a resolution of 512 $\times$ 512 pixels. All synthetic data and code are available on our project page (see Abstract).
@@ 4.1.2 Raw Telescope Image Pre-processing
We first validate our pre-processing pipeline on the simulated data. Following our pipeline—lucky imaging and stacking (every 100 images), wavelet-based multi-scale sharpening, and VRT denoising—we reduce the 14,000 distorted frames to 140 high-quality images per satellite model. We compared our method against other cutting-edge deblurring algorithms. Table I summarizes the evaluation metrics (SSIM, PSNR, LPIPS) across various preprocessing methods. The results indicate that our proposed pipeline (Stacking + Sharpening + VRT) consistently surpasses pure deep learning approaches (e.g., Restormer[55]), traditional techniques (e.g., Stacking + Sharpening), self-supervised methods, and other hybrid strategies, specifically in PSNR and SSIM metrics.
@@ 4.1.3 Robust Pose Initialization
From the 140 processed images, we select every tenth frame (e.g., 1st, 10th, 20th) for the training set and reserve the intermediate frames for validating the reconstruction. We first compare the pose estimation accuracy of our method against existing techniques, including feature-based SfM methods—such as Colmap[31], hierarchical localization[56], and detector-free SfM[57]—as well as several feed-forward deep learning approaches, including DUSt3R[58] and RayDiffusion[59]. Fig. 5 and Table II compare the estimated poses to the ground truth (GT) on the training set. Feature-based SfM methods accumulate noise, which leads to inaccurate keypoint matching and causes bundle adjustment (BA) to fail to converge; consequently, no reasonable pose estimation can be reported. Similarly, deep learning methods struggle with extreme focal lengths, structural symmetry, and poor textures, resulting in incorrect depth estimation and subsequent pose errors. Other approaches, such as PoseDiffusion[60], MV-DUSt3R+[61], and Colmap-free 3DGS[62], also exhibit substantial pose estimation errors. In contrast, our pose initialization method achieves an ATE of only 2.61, with the estimated pose distribution closely aligning with the ground truth. As explained in Sec.3.2.1, our superior results stem from three key factors: (1) a robust orthographic projection model; (2) the use of RANSAC algorithms to eliminate erroneous matches; and (3) a strategy that leverages only the matching relationships between adjacent frames to relax the convergence conditions. These design choices enable our method to overcome the limitations of traditional SfM and deep learning approaches, significantly enhancing reliability for pose estimation of texture-poor space objects.